### v1

# TempFlow-GRPO: When Timing Matters for GRPO in Flow Models

Abstract
Recent flow matching models for text-to-image generation have achieved remarkable quality, yet their integration with reinforcement learning for human preference alignment remains suboptimal, hindering fine-grained reward-based optimization. We observe that the key impediment to effective GRPO training of flow models is the temporal uniformity assumption in existing approaches: sparse terminal rewards with uniform credit assignment fail to capture the varying criticality of decisions across generation timesteps, resulting in inefficient exploration and suboptimal convergence. To remedy this shortcoming, we introduce TempFlow-GRPO (Temporal Flow GRPO), a principled GRPO framework that captures and exploits the temporal structure inherent in flow-based generation. TempFlow-GRPO introduces two key innovations: (i) a trajectory branching mechanism that provides process rewards by concentrating stochasticity at designated branching points, enabling precise credit assignment without requiring specialized intermediate reward models; and (ii) a noise-aware weighting scheme that modulates policy optimization according to the intrinsic exploration potential of each timestep, prioritizing learning during high-impact early stages while ensuring stable refinement in later phases. These innovations endow the model with temporally-aware optimization that respects the underlying generative dynamics, leading to state-of-the-art performance in human preference alignment and standard text-to-image benchmarks.
1 Introduction
While text-to-image diffusion models have achieved unprecedented visual quality and semantic control Esser et al. (2024); Xie et al. (2024a); Labs et al. (2025), aligning their outputs with human preference remains a formidable challenge. Reinforcement learning has emerged as a promising solution, giving rise to the field of Diffusion RL Wallace et al. (2024); Black et al. (2023); Fan et al. (2023). However, the performance of these methods remains suboptimal, hindered by two fundamental limitations that have been largely overlooked: ignoring the temporal dynamics of generation and lacking intermediate feedback signals. These approaches apply uniform optimization across all timesteps and provide rewards only at completion, missing the varying importance of decisions throughout the generation process.
The majority of existing approaches Gu et al. (2024); Hong et al. (2024), including recent works like Flow-GRPO Liu et al. (2025) and DanceGRPO Xue et al. (2025), treat the multi-step generation process as a ”black box” with temporally agnostic optimization. They apply uniform updates across all timesteps despite the fact that each timestep operates under different noise conditions and contributes differently to final image quality. Specifically, we plot the Fig. 1 (left) with applying SDE at only one timestep in the entire ODE trajectory, which ensures that any deviation in the final reward can be attributed to the stochastic exploration introduced at that specific step. As shown in Fig. 1 (left), the std of the reward varies dramatically between timesteps, reaching a peak during early structural decisions (steps 0-2) and approaching zero during final refinements (steps 6-8). Yet Flow-GRPO maintains uniform treatment throughout, squandering high-impact exploration opportunities while focusing on later steps. Alternative approaches like SPO Liang et al. (2025) attempt to address temporal dynamics through process reward models, but training such models on semantically ambiguous intermediate states is notoriously difficult. This raises a fundamental question: how can we effectively achieve precise credit assignment for intermediate actions while adapting optimization intensity to each timestep’s exploration capacity?
Figure 1: (Left) Reward Variance Analysis: We plot the standard deviation of PickScore at each denoising step for 200 prompts, per prompt group size is 24. The results, obtained via applying SDE at only one step, reveal that reward variance is highest in the initial steps, indicating that early-stage interventions are most impactful for exploration. (Right) Method Illustration: By branching a stochastic (SDE) exploration from a specific, known state on a deterministic (ODE) trajectory, we create a controlled experiment. The resulting difference in the final reward can be unambiguously attributed to the exploration action taken at that precise branching point.
We address these limitations with TempFlow-GRPO, a temporally-aware RL framework built on two key insights. First, we define the visualization method in the left of Fig. 1 as trajectory branching, which enables precise credit assignment by strategically introducing stochasticity at individual timesteps while maintaining deterministic evolution elsewhere (Fig. 1, Right). This provides provable guarantees: (1) reward variance localizes to the branching point, (2) improvements are directly attributable to specific exploration outcomes, and (3) existing reward models require no modification. Second, noise-aware policy weighting modulates optimization intensity based on each timestep’s intrinsic noise level. Early high-noise stages receive larger weight updates to encourage structural exploration, while late low-noise stages receive gentler updates to preserve learned features. Together, these mechanisms create a framework that is conceptually simple, computationally efficient, and seamlessly integrates into existing flow matching architectures—all while respecting the temporal dynamics that uniform approaches ignore.
Our main contributions are threefold:
  • We pinpoint temporal uniformity—the equal treatment of all timesteps—as the primary limitation of flow-based GRPO. Our proposed TempFlow-GRPO overcomes this by introducing two key innovations: precise credit assignment to intermediate actions and noise-aware adaptation of optimization intensity.
  • We introduce trajectory branching and noise-aware weighting to learn temporally-structured policies that respect generative dynamics.
  • We demonstrate state-of-the-art performance on standard text-to-image benchmarks, achieving superior sample quality, human preference alignment, and compositional image generation compared to existing flow-based RL methods.
2 Related Work
Alignment for Diffusion Models. Alignment for diffusion models has become a rapidly emerging topic in generative modeling research. D3PO Yang et al. (2024) introduce the Direct Preference for Denoising Diffusion Policy Optimization method to directly fine-tune diffusion models. Diffusion-DPO Wallace et al. (2024) is adapted from the Direct Preference Optimization (DPO Rafailov et al. (2023)), a simpler alternative to RLHF which directly optimizes a policy that best satisfies human preferences under a classification objective. DyMO Xie & Gong (2025) propose a plug-and-play training-free alignment method for aligning the generated images and human preferences during inference. Recently, Flow-GRPO Liu et al. (2025), the first method to integrate online reinforcement learning (RL) into flow matching models. However, Flow-GRPO applies uniform optimization pressure across all timesteps and suffers from sparse terminal reward problems, failing to account for the time-varying exploration potential inherent in the stochastic diffusion process. Our TempFlow-GRPO addresses these limitations through trajectory branching for precise credit assignment and noise-aware policy weighting that aligns optimization pressure with the natural exploration capacity at each timestep.
Figure 2: (Left) Performance comparison on the PickScore benchmark. To ensure a fair and robust evaluation, we first present Flow-GRPO (Fixed), an improved baseline where we replace the original global std stabilization with a more appropriate group-wise std. On top of this stronger baseline, our core innovation, trajectory branching (+ Traj. Branch), delivers a significant performance leap. Our full model, TempFlow-GRPO, integrates all components to achieve the highest performance. (Right) TempFlow-GRPO’s dominance is further confirmed on the Geneval benchmark, where it substantially surpasses not only the Flow-GRPO baseline but also leading state-of-the-art models, including GPT-4o, SD3.5-M, and FLUX.
Process Reward. Recent studies have demonstrated that shaping the reward process, rather than solely relying on sparse terminal rewards, can significantly accelerate learning and improve policy performance. Zhang et al. (2025) introduces a more comprehensive evaluation framework, which combines response-level and step-level metrics. ThinkPRM Khalifa et al. (2025) builds data-efficient PRMs as verbalized step-wise reward models that verify every step in the solution by generating a verification chain-of-thought (CoT). In diffusion models, SPO Liang et al. (2024) trains a separate step-aware preference model that can be applied to both noisy and clean images. Despite this, PRMs require step-level supervision, making them expensive to train. There are several methods using outcome reward to replace process reward. PRIME Cui et al. (2025) enables online PRM updates using only policy rollouts and outcome labels through implict process rewards. Math-Shepherd Wang et al. (2023) assigns a reward score to each step of math problem solutions. Due to the significant challenges in scoring noisy images, there is an urgent need for an algorithm that enables process reward on flow models. Our TempFlow-GRPO elegantly circumvents the need for specialized process reward models by directly attributing outcome-based reward signals to intermediate exploratory actions, enabling precise credit assignment without the computational overhead of training step-level evaluators for semantically ambiguous noisy states.
3 Preliminary: Flow-GRPO
Flow-GRPO enhances flow models using online RL. First, we revisit the core idea of GRPO. Then, we show how Flow-GRPO converts the deterministic ODE sampler into a SDE sampler with same marginal distribution, which introduces the stochasticity needed for applying GRPO.
GRPO. RL aims to learn a policy that maximizes the expected cumulative reward. GRPO optimizes the policy model by maximizing the following objective:
$\mathcal{J}_{\text{Flow-GRPO}}(\theta)=\mathbb{E}_{\bm{c}\sim\mathcal{C},\{\bm{x}^{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|\bm{c})}f(r,\hat{A},\theta,\epsilon,\beta)$ (1)
where
$\begin{split}f(r,\hat{A},\theta,\epsilon,\beta)&=\frac{1}{G}\sum\limits_{i=1}^{G}\frac{1}{T}\sum\limits_{t=0}^{T-1}(\mathop{min}(r_{t}^{i}(\theta)\hat{A}_{t}^{i},\text{clip}(r_{t}^{i}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}^{i})-\beta D_{KL}(\pi_{\theta}||\pi_{\text{ref}}))\\ r_{t}^{i}(\theta)&=\frac{p_{\theta}(\bm{x}^{i}_{t-1}|\bm{x}_{t}^{i},\bm{c})}{p_{\theta_{\text{old}}}(\bm{x}^{i}_{t-1}|\bm{x}_{t}^{i},\bm{c})},T\text{ is the timestep.}\end{split}$ (2)
Given a prompt $\bm{c}$, the flow model $p_{\theta}$ samples a group of $G$ individual images $\{\bm{x}_{0}^{i}\}_{i=1}^{G}$ and the corresponding reverse-time trajectories $\{(\bm{x}_{T}^{i},\bm{x}_{T-1}^{i},...,\bm{x}_{0}^{i}\}_{i=1}^{G}$. Then, the advantage of the $i$-th image is calculated by normalizing the group-level rewards as follows:
$\hat{A}_{t}^{i}=\frac{R(\bm{x}_{0}^{i},\bm{c})-\text{mean}(\{R(\bm{x}^{i}_{0},\bm{c})\}_{i=1}^{G})}{\text{std}(\{R(\bm{x}_{0}^{i},\bm{c})\}_{i=1}^{G})}$ (3)
Convert ODE to SDE. GRPO relies on stochastic sampling to generate diverse trajectories for advantage estimation and exploration. However, flow matching models use a deterministic ODE for the forward process:
$d\bm{x}_{t}=\bm{v}_{t}dt$ (4)
Flow-GRPO converts the deterministic ODE into an equivalent SDE that matches the original model’s marginal probability density function at all timesteps. The final update rule is as follows:
$\bm{x}_{t+\Delta t}=\bm{x}_{t}+[\bm{v}_{\theta}(\bm{x}_{t},t)+\frac{\sigma_{t}^{2}}{2t}(\bm{x}_{t}+(1-t)\bm{v}_{\theta}(\bm{x}_{t},t))]\Delta t+\sigma_{t}\sqrt{\Delta t}\bm{\epsilon}$ (5)
where $\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})$ injects stochasticity and $\sigma_{t}=a\sqrt{\frac{t}{1-t}}$. And the KL divergence between $\pi_{\theta}$ and the reference policy $\pi_{\text{ref}}$ is a closed form:
[…]
Figure 3: Overview of TempFlow-GRPO Framework. Our method performs trajectory branching by switching from ODE to SDE sampling at selected timesteps (t=k, j, i), injecting noise $\sigma_{t}\sqrt{\Delta t}\epsilon$ to create exploratory branches. Each branch generates a distinct outcome with reward $R_{i}$, enabling precise credit assignment. The framework applies noise-aware weighting where $\omega_{i}>\omega_{j}>\omega_{k}$, prioritizing optimization at high-noise early stages (larger circles) over low-noise refinement phases (smaller circles), aligning learning intensity with each timestep’s intrinsic exploration capacity. We visualize the model’s learning process as an astronaut exploring unknown planets: in early stages , the model explores vast possibility spaces with high uncertainty, while later stages involve focused navigation toward the final destination.
4 Methods
4.1 Temporal Flow-GRPO
While Flow-GRPO represents a significant advancement in applying online RL to flow matching models, its direct adoption of the standard GRPO framework overlooks the unique, time-dependent dynamics inherent in the generative process. We identify two fundamental limitations that motivate our approach:
Sparse Terminal Reward Problem: The current approach relies on sparse terminal rewards, assigning the same credit uniformly across all timesteps. This fails to differentiate the critical, high-impact decisions of early generation from the fine-tuning adjustments of later stages.
Uniform Optimization Weighting: The GRPO objective applies uniform optimization among all steps, ignoring that the underlying SDE sampler possesses non-uniform noise. The potential for meaningful exploration is greater in high-noise early steps than in the low-noise refinement phase Xie & Gong (2025).
To address these limitations, as illustrated in Fig. 3, we propose TempFlow-GRPO, a unified framework that introduces two synergistic innovations: process rewards via trajectory branching and noise-aware policy weighting.
4.1.1 Trajectory Branching for Process Rewards
Traditional process reward methods require specialized reward models to evaluate noisy intermediate states $\bm{x}_{t}$, which is exceptionally difficult due to the semantic ambiguity of partially-denoised representations. We propose an elegant alternative that leverages the deterministic-stochastic sampling methods of flow matching models.
Key Insight: Instead of training complex process reward models, we use existing high-quality outcome-based reward models and directly attribute their scores to intermediate exploratory actions through a novel trajectory branching mechanism.
Consider a generative process parameterized by $\theta$. In flow matching models, the policy gradient can be written as:
$\nabla_{\theta}\mathcal{J}(\theta)=\sum_{k=0}^{T-1}\mathbb{E}_{\bm{x}_{T}\sim\mathcal{N}(0,\bm{I}),\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})}[\nabla_{\theta}\log p_{\theta}(\bm{x}_{k-1}|\bm{x}_{k})\hat{A}_{k}]$ (7)
Definition 1 (Trajectory Branching): We define a trajectory branching operation where a trajectory evolves deterministically at a designated branching timestep $k$, where $\bm{x}_{k}$ is sampled from initial noise $\bm{x}_{T}$ using Eq. 4. At this branching point, stochasticity is introduced via noise variable $\bm{\epsilon}$ in Eq. 5, yielding $\bm{x}_{k-1}=\text{SDE}(\bm{x}_{k},\bm{\epsilon})$. The remainder of the trajectory $\bm{x}_{k-2},\bm{x}_{k-3},\ldots,\bm{x}_{0}$ is generated deterministically as $\bm{x}_{0}=\text{ODE}(\bm{x}_{k-1})$.
Theorem 1 (Credit Localization): Since all stochasticity and model controllability are concentrated at the branching point, the total reward variance and all parameter-dependent improvements are entirely attributable to the outcome of noise injection at $k$. This enables rigorous and efficient credit assignment localized to the branching point.
In practice, we replace the reward for the $k$-th step from $R(\bm{x}_{0}^{i},\bm{c})$ to $R(\text{ODE}(\text{SDE}(\bm{x}_{k}^{i},\bm{\epsilon}^{i})),\bm{c})$, where $\bm{x}_{0}^{i}$ is sampled with SDE and $\text{ODE}(\text{SDE}(\bm{x}_{k}^{i},\bm{\epsilon}^{i})$ is sampled with ODE-SDE-ODE. Trajectory branching allows for the precise attribution of the terminal reward to step $k$, effectively creating a temporally-aware reward signal.
Figure 4: (Left) A comparative analysis between the reward standard deviation (Reward Std) and the noise level over the generative steps. The two curves show a strong correlation. (Right) Scale terms reveal a mismatch in standard GRPO: Scale terms are inversely proportional to noise level, causing low-noise refinement steps to dominate optimization despite having minimal impact on image content. Early steps that establish global structure receive weak gradients, while late steps that only adjust local details produce strong gradients. Our noise-aware reweighting compensates for this inverse relationship, ensuring that optimization intensity aligns with each timestep’s actual capacity to influence the final image.
4.1.2 Noise-Aware Policy Weighting
While trajectory branching provides precise reward signals for individual timesteps, the generative process consists of a sequence of $T$ potential branching points with fundamentally different characteristics. The SDE sampler exhibits time-varying stochasticity: the noise injection magnitude $\sigma_{t}\sqrt{\Delta t}$ is large during initial generation stages and diminishes to near zero during final refinement stages. This non-uniform noise distribution implies that exploration capacity varies dramatically across timesteps. An exploratory action at an early stage has vastly different impact and risk compared to perturbations on near-perfect images. However, standard GRPO applies uniform optimization pressure, implicitly assuming equal learning importance at all stages.
Reweighting by Reward Std. An intuitive approach to reweighting the policy loss is to utilize the standard deviation of the rewards. As illustrated in the left of Fig. 4, the reward standard deviation exhibits a similar decaying trend as the denoising process unfolds, suggesting it could serve as a meaningful indicator of learning importance. However, this approach introduces significant complexity. It would necessitate maintaining a dynamic weighting hyperparameter for each timestep, likely tracked via an Exponential Moving Average (EMA Morales-Brotons et al. (2024)). This dynamic estimation process can be unstable and adds an undesirable layer of complexity to the training regime. Consequently, we seek a more direct and elegant weighting scheme.
Reweighting by Noise Level. We wonder if directly leverage the level of the exploration space itself as a proxy for a reweighting factor. We visualized the noise level alongside the reward standard deviation (Fig. 4, left) and observed a striking correspondence between the two. This strong correlation suggests that the noise level, serves as an excellent and intrinsic proxy for the exploration capacity and associated risk at each timestep.
We therefore propose to reweight the policy loss directly using the noise level. For each timestep $t$, we introduce a weighting factor proportional to the noise level. Specifically, we modify the original policy gradient loss function to the following weighted form:
$\mathcal{J}_{\text{policy}}(\theta)=\frac{1}{G}\sum\limits_{i=1}^{G}\frac{1}{T}\sum\limits_{t=0}^{T-1}\text{Norm}(\sigma_{t}\sqrt{\Delta t})(\mathop{min}(r_{t}^{i}(\theta)\hat{A}_{t}^{i},\text{clip}(r_{t}^{i}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}^{i})$ (8)
The intuition behind this weighting strategy is to align the optimization pressure with the inherent properties of the generative process. In the early stages of generation, noise is large, amplifying the learning signal during these high-noise, high-impact phases and encouraging the model to perform effective macroscopic exploration. As generation proceeds, noise diminishes, shifts the optimization focus towards fine-grained adjustments and stability, preventing aggressive exploration from corrupting a high-fidelity state with noise or artifacts.
4.2 Policy Gradient-Based Theoretical Justification
To provide a deeper understanding of our approach, we now examine it from the policy gradient perspective. Notice that in the $\sum_{k=0}^{T-1}$, k is the timestep, and in the equation, k is the value of the timestep. Simplifying Eq. 5, we obtain $\bm{x}_{k-1}\sim\mathcal{N}(\mu_{\theta}(\bm{x}_{k},k),\sigma_{k}^{2}\Delta k\bm{I})$, where:
$\mu_{\theta}(\bm{x}_{k},k)=\bm{x}_{k}+\left[\bm{v}_{\theta}(\bm{x}_{k},k)+\frac{\sigma_{k}^{2}}{2k}\left(\bm{x}_{k}+(1-k)\bm{v}_{\theta}(\bm{x}_{k},k)\right)\right]\Delta k$ (9)
Starting from the policy gradient formulation in Eq. 7, we have:
$\nabla_{\theta}\mathcal{J}(\theta)=\sum_{k=0}^{T-1}\mathbb{E}_{\bm{x}_{T}\sim\mathcal{N}(0,\bm{I}),\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})}[\nabla_{\theta}\log p_{\theta}(\bm{x}_{k-1}|\bm{x}_{k})\hat{A}_{k}]$ (10)
Substituting $\bm{x}_{k-1}$ in the log-probability:
$\nabla_{\theta}\mathcal{J}(\theta)$ $=\sum_{k=0}^{T-1}\mathbb{E}_{\bm{x}_{T}\sim\mathcal{N}(0,\bm{I}),\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})}\left[\nabla_{\theta}\log\exp\left(-\frac{\|\bm{x}_{k-1}-\mu_{\theta}(\bm{x}_{k},k)\|^{2}}{2\sigma_{k}^{2}\Delta k}\right)\hat{A}_{k}\right]$ (11)
$=\sum_{k=0}^{T-1}\mathbb{E}_{\bm{x}_{T}\sim\mathcal{N}(0,\bm{I}),\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})}\left[\nabla_{\theta}\left(-\frac{\|\bm{x}_{k-1}-\mu_{\theta}(\bm{x}_{k},k)\|^{2}}{2\sigma_{k}^{2}\Delta k}\right)\hat{A}_{k}\right]$ (12)
Taking the gradient with respect to $\theta$:
$\nabla_{\theta}\mathcal{J}(\theta)=\sum_{k=0}^{T-1}\mathbb{E}_{\bm{x}_{T}\sim\mathcal{N}(0,\bm{I}),\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})}\left[\frac{\bm{x}_{k-1}-\mu_{\theta}(\bm{x}_{k},k)}{\sigma_{k}^{2}\Delta k}\cdot\nabla_{\theta}\mu_{\theta}(\bm{x}_{k},k)\hat{A}_{k}\right]$ (13)
Since $\bm{x}_{k-1}=\mu_{\theta}(\bm{x}_{k},k)+\sigma_{k}\sqrt{\Delta k}\cdot\bm{\epsilon}$ where $\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})$:
$\nabla_{\theta}\mathcal{J}(\theta)=\sum_{k=0}^{T-1}\mathbb{E}_{\bm{x}_{T}\sim\mathcal{N}(0,\bm{I}),\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})}\left[\frac{\bm{\epsilon}}{\sigma_{k}\sqrt{\Delta k}}\cdot\nabla_{\theta}\mu_{\theta}(\bm{x}_{k},k)\hat{A}_{k}\right]$ (14)
Expanding $\nabla_{\theta}\mu_{\theta}(\bm{x}_{k},k)$:
$\nabla_{\theta}\mu_{\theta}(\bm{x}_{k},k)$ $=\nabla_{\theta}\left[\bm{x}_{k}+\left(\bm{v}_{\theta}(\bm{x}_{k},k)+\frac{\sigma_{k}^{2}}{2k}\left(\bm{x}_{k}+(1-k)\bm{v}_{\theta}(\bm{x}_{k},k)\right)\right)\Delta k\right]$ (15)
$=\nabla_{\theta}\left[\left(\bm{v}_{\theta}(\bm{x}_{k},k)+\frac{\sigma_{k}^{2}(1-k)}{2k}\bm{v}_{\theta}(\bm{x}_{k},k)\right)\Delta k\right]$ (16)
$=\left(1+\frac{\sigma_{k}^{2}(1-k)}{2k}\right)\Delta k\cdot\nabla_{\theta}\bm{v}_{\theta}(\bm{x}_{k},k)$ (17)
Substituting back:
$\nabla_{\theta}\mathcal{J}(\theta)$ $=\sum_{k=0}^{T-1}\mathbb{E}_{\bm{x}_{T}\sim\mathcal{N}(0,\bm{I}),\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})}\left[\frac{\bm{\epsilon}}{\sigma_{k}\sqrt{\Delta k}}\cdot\left(1+\frac{\sigma_{k}^{2}(1-k)}{2k}\right)\Delta k\cdot\nabla_{\theta}\bm{v}_{\theta}(\bm{x}_{k},k)\hat{A}_{k}\right]$ (18)
$=\sum_{k=0}^{T-1}\mathbb{E}_{\bm{x}_{T}\sim\mathcal{N}(0,\bm{I}),\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})}\left[\left(\frac{\sqrt{\Delta k}}{\sigma_{k}}+\frac{\sigma_{k}\sqrt{\Delta k}(1-k)}{2k}\right)\cdot\bm{\epsilon}\cdot\nabla_{\theta}\bm{v}_{\theta}(\bm{x}_{k},k)\hat{A}_{k}\right]$ (19)
With $\sigma_{k}=a\sqrt{\frac{k}{1-k}}$, we get:
$\frac{\sqrt{\Delta k}}{\sigma_{k}}$ $=\frac{\sqrt{\Delta k}}{a\sqrt{\frac{k}{1-k}}}=\frac{1}{a}\sqrt{\frac{\Delta k(1-k)}{k}}$ (20)
$\frac{\sigma_{k}\sqrt{\Delta k}(1-k)}{2k}$ $=\frac{a\sqrt{\frac{k}{1-k}}\sqrt{\Delta k}(1-k)}{2k}=\frac{a}{2}\sqrt{\frac{\Delta k(1-k)}{k}}$ (21)
Therefore:
$\nabla_{\theta}\mathcal{J}(\theta)=\sum_{k=0}^{T-1}\mathbb{E}_{\bm{x}_{T}\sim\mathcal{N}(0,\bm{I}),\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})}\left[\left(\frac{1}{a}+\frac{a}{2}\right)\underbrace{\sqrt{\frac{\Delta k(1-k)}{k}}}_{\text{Scale Term}}\cdot\bm{\epsilon}\cdot\nabla_{\theta}\bm{v}_{\theta}(\bm{x}_{k},k)\hat{A}_{k}\right]$ (22)
This reveals that the natural gradient coefficient is proportional to $\sqrt{\frac{1-k}{k}}\sqrt{\Delta k}$, which captures the intrinsic exploration potential at timestep $k$. After reweighting, we have the following derivation:
$\nabla_{\theta}\mathcal{J}(\theta)=\sum_{k=0}^{T-1}\mathbb{E}_{\bm{x}_{T}\sim\mathcal{N}(0,\bm{I}),\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})}\left[\left(\frac{1}{a}+\frac{a}{2}\right)\underbrace{\Delta k}_{\text{Scale Term}}\cdot\bm{\epsilon}\cdot\nabla_{\theta}\bm{v}_{\theta}(\bm{x}_{k},k)\hat{A}_{k}\right]$ (23)
Consider $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})}[\bm{\epsilon}\hat{A}_{k}]$, suppose the final reward is a function of the small noise vector $\sigma_{k}\sqrt{\Delta k}\bm{\epsilon}$ applied at a certain step. When $\sigma_{k}\sqrt{\Delta k}\bm{\epsilon}$ is small (and drawn from a zero-mean Gaussian), we can approximate the reward using a first-order Taylor expansion:
$R_{k}(\sigma_{k}\sqrt{\Delta k}\bm{\epsilon})\approx R_{k}(0)+\sigma_{k}\sqrt{\Delta k}\bm{\epsilon}^{T}\nabla_{\sigma_{k}\sqrt{\Delta k}\bm{\epsilon}}R_{k}|_{\sigma_{k}\sqrt{\Delta k}\bm{\epsilon}=0}$ (24)
Since $\hat{A}_{k}$ is normalized version of $R_{k}$, the mean and std are as follows (let $\nabla_{\sigma_{k}\sqrt{\Delta k}\bm{\epsilon}}R_{k}|_{\sigma_{k}\sqrt{\Delta k}\bm{\epsilon}=0}=g_{k}$):
[…]
Therefore:
[…]
Eq. 26 indicates that the norm of $\mathbb{E}_{\bm{\epsilon}}[\bm{\epsilon}\hat{A}_{k}]$ is invariant among the timesteps. So in Eq. 22 and Eq. 23, the scale terms that modulate the contribution of each timestep’s model gradient $\nabla_{\theta}\bm{v}_{\theta}(\bm{x}_{k},k)$ to the overall reward gradient simplify to distinct functions, which we denote as $\sqrt{\frac{\Delta k(1-k)}{k}}$ and $\Delta k$. We visualize these scale terms under different flow shift, as depicted in the right of Fig. 4. As our analysis indicates, the setting of “w/o reweighting” exhibits a highly imbalanced contribution from each timestep’s model gradient to the final reward gradient. This non-uniformity systematically causes the contribution from the early denoising step where the model performs broad, structural exploration to be significantly smaller than the contribution from the late step, which focuses on fine-grained refinement. By employing our proposed noise-aware policy reweighting, this issue is substantially mitigated, as the scale term simplifies to being directly proportional to the step size, $\Delta k$. Furthermore, when the flow shift is set as 1, our method achieves a perfect equilibrium: it ensures that the gradient contribution from every single timestep is precisely equal, thereby completely balancing this effect across the entire generation trajectory.
5 Experiment
Following Flow-GRPO, we validate our approach on Compositional Image Generation in Geneval Ghosh et al. (2023) and Human Preference Alignment in PickScore Kirstain et al. (2023). To ensure a fair comparison, we normalized the weights applied to the policy loss to have a mean of 1 at all timesteps. We use num groups as 48 and group size as 24. The beta of kl loss, is set to 0.001 for PickScore and 0.004 in Geneval. Unless otherwise specified, the image size is set to 512, consistent with the configuration used for Flow-GRPO. It is worth noting that we further improve upon the original Flow-GRPO method by proposing a variant, Flow-GRPO (FIXED), which demonstrates superior performance. In our PickScore experiments, we adopt this improved version to ensure a fair and strong baseline for comparison.
5.1 Main Results
Table 1: GenEval Result. Best scores are in blue, second-best in green. Results for models are from Flow-GRPO. Obj.: Object; Attr.: Attribution.
Model Step Overall Single Obj. Two Obj. Counting Colors Position Attr. Binding
Diffusion Models
LDM Rombach et al. (2022) - 0.37 0.92 0.29 0.23 0.70 0.02 0.05
SD1.5 Rombach et al. (2022) - 0.43 0.97 0.38 0.35 0.76 0.04 0.06
SD2.1 Rombach et al. (2022) - 0.50 0.98 0.51 0.44 0.85 0.07 0.17
SD-XL Podell et al. (2023) - 0.55 0.98 0.74 0.39 0.85 0.15 0.23
DALLE-2 Ramesh et al. (2022) - 0.52 0.94 0.66 0.49 0.77 0.10 0.19
DALLE-3 Betker et al. (2023) - 0.67 0.96 0.87 0.47 0.83 0.43 0.45
Autoregressive Models
Show-o Xie et al. (2024b) - 0.53 0.95 0.52 0.49 0.82 0.11 0.28
Emu3-Gen Wang et al. (2024) - 0.54 0.98 0.71 0.34 0.81 0.17 0.21
JanusFlow Ma et al. (2025) - 0.63 0.97 0.59 0.45 0.83 0.53 0.42
Janus-Pro-7B Chen et al. (2025) - 0.80 0.99 0.89 0.59 0.90 0.79 0.66
GPT-4o Hurst et al. (2024) - 0.84 0.99 0.92 0.85 0.92 0.75 0.61
Flow Matching Models
FLUX.1 Dev Black et al. (2025) - 0.66 0.98 0.81 0.74 0.79 0.22 0.45
SD3.5-L Esser et al. (2024) - 0.71 0.98 0.89 0.73 0.83 0.34 0.47
SANA-1.5 4.8B Xie et al. (2025) - 0.81 0.99 0.93 0.86 0.84 0.59 0.65
SD3.5-M Esser et al. (2024) - 0.63 0.98 0.78 0.50 0.81 0.24 0.52
GRPO based Methods
SD3.5-M+Flow-GRPO Liu et al. (2025) 5600 0.95 1.00 0.99 0.95 0.92 0.99 0.86
SD3.5-M+Flow-GRPO Liu et al. (2025) 4400 0.90 0.99 0.97 0.90 0.88 0.85 0.80
SD3.5-M+TempFlow-GRPO 4400 0.97 1.00 0.99 0.96 0.97 0.98 0.90
Compositional Image Generation. We evaluate the compositional image generation capability of TempFlow-GRPO on the Geneval benchmark with its corresponding reward model. The experimental results are summarized in Tab. 1. As shown, our approach significantly improves the performance of the base model, increasing the overall score from 0.63 to 0.97. Furthermore, among GRPO-based methods, our method substantially outperforms Flow-GRPO: it achieves a performance of 0.97 within only 4,400 steps, whereas Flow-GRPO reaches only 0.90 under the same conditions. Additionally, as illustrated in Fig. 2, our method requires only about 2,000 steps to achieve a score of 0.95, while Flow-GRPO needs approximately 5,600 steps to reach the same level. Overall, these results demonstrate that TempFlow-GRPO not only accelerates convergence but also achieves superior final performance compared to existing approaches.
Human Preference Alignment. To further validate the generalizability of our approach, we conducted experiments on the PickScore benchmark, using PickScore as the reward model. As shown in the left of Fig. 2, our method, TempFlow-GRPO, achieves the highest performance, surpassing the original Flow-GRPO by approximately 1.3% and outperforming the improved baseline Flow-GRPO (Fixed) by about 0.6 %. Notably, our method requires only 100–200 training steps to match the performance of Flow-GRPO, and just 300–400 steps to reach the level of Flow-GRPO (Fixed). These results on PickScore further demonstrate the general applicability of our method as a unified flow-based RL algorithm across different reward models.
5.2 Analysis
Figure 5: Ablation Studies on trajectory branch and noise-aware policy reweighting. Left: PickScore Benchmark. Right: Geneval Benchmark.
Ablation of TempFlow-GRPO. We conducted ablation studies to explore the effectiveness of our proposed components: the trajectory branch and noise-aware policy weighting. These ablations were performed on both the Geneval and PickScore benchmarks. As shown in the Fig. 5, on the PickScore benchmark, introducing the trajectory branch further improves the performance of Flow-GRPO (Fixed), and applying noise-aware reweighting on top of this yields the highest pickscore results. On the Geneval benchmark, the benefit of the noise-aware strategy is even more significant: compared to Flow-GRPO, noise-aware policy reweighting boosts performance from 0.85 to 0.94, a 9% improvement while the trajectory branch also brings about a substantial gain of approximately 5%. These ablation results clearly demonstrate the effectiveness of our proposed methods.
Figure 6: Comparsion on PickScore benchmark in 1024 resolution.
Result of 1024 resolution. We further explored the effectiveness of our approach across different image resolutions, using PickScore as the reward model. As illustrated in the Fig. 6, after 450 training steps, TempFlow-GRPO achieves a 0.6% improvement in PickScore. Notably, our method requires only about 180 steps to reach the performance that Flow-GRPO (Fixed) attains after 450 steps. This further demonstrates the efficiency and effectiveness of our proposed method across varying resolutions.
Qualitative Result. We also conducted qualitative analyses on the SD3.5-M, Flow-GRPO (Fixed), and TempFlow-GRPO. As shown in the visualization results in Fig 7, compared to Flow-GRPO (Fixed), TempFlow-GRPO produces images with noticeably finer details and fewer visual artifacts or mistakes. In particular, our approach demonstrates superior capability in preserving complex structures and realistic textures. These qualitative improvements further highlighting the advantages of our method in generating high-quality, visually appealing images.
Figure 7: Qualitative comparison between SD3.5-M, Flow-GRPO and TempFlow-GRPO with PickScore as reward. We use the red box to indicate the error region.
6 Limitation and Future Work
Although our method achieves significant improvements in both performance and image quality, the current experiments are based on a single reward model. In future work, we plan to focus on incorporating multi rewards from more powerful models, aiming to enhance performance across multiple dimensions. Additionally, we intend to design a comprehensive pipeline that can systematically improve various aspects of generative quality.
7 Conclusion
We presented TempFlow-GRPO, a temporally-aware reinforcement learning framework that addresses fundamental limitations in existing flow-based GRPO methods. Our key insight is that the uniform treatment of all timesteps creates a misalignment between optimization effort and actual impact on generation quality. Through trajectory branching, we enable precise credit assignment to intermediate actions without requiring specialized process reward models. Through noise-aware weighting, we ensure that optimization intensity matches each timestep’s exploration potential, preventing both under-exploration of critical early decisions and over-optimization of minor refinements. Our extensive experiments demonstrate that TempFlow-GRPO achieves state-of-the-art performance on human preference alignment and compositional generation benchmarks, surpassing both traditional GRPO methods and recent flow-based approaches.